The average American city has 50 buses that use green energy, with a standard deviation of 5 buses. A city planner is interested in the sample mean number of buses for a sample of 30 cities. He calculates that: μx=μn−−√=5030−−√=9.13 and σx¯=σn−−√=530−−√=0.91 What did he do wrong?
Answer #1

\mu_{\bar x} = \mu = 50

\sigma_{\bar x} = \sigma/\sqrt n = 5/V30 = 0.9129